OpenStudy (aakashsudhakar):

Integral here, taken from my Bioengineering Transport Phenomena class regarding heat transfer out of a particle. See below if you dare!

OpenStudy (aakashsudhakar):

\[T*e ^{\frac{ 2br ^{2} - ar ^{4} }{ 4 }} = C _{1} \int\limits_{ }^{ } \frac{ 1 }{ r } e ^{\frac{ 2br ^{2} - ar ^{4} }{ 4 }} dr\]

OpenStudy (aakashsudhakar):

T = T(r) and thus is a function of r, so it makes it very difficult or practically impossible to derive both sides and attempt to find the derivative of the left side.

OpenStudy (jerry45):

wow i feel bad for you

OpenStudy (irishboy123):

why not just thrash it out?! there's a good chance it'll separate for ya. a lot of stuff is just gonna cancel out.

OpenStudy (irishboy123):

But it looks like it is in integrating factor form already, so you may just go round in a circle :-/